### v1

# After Retrieval, Before Generation: Enhancing the␣Trustworthiness of Large Language Models in RAG

Abstract
Retrieval-augmented generation (RAG) systems face critical challenges in balancing internal (parametric) and external (retrieved) knowledge, especially when these sources conflict or are unreliable. To analyze these scenarios comprehensively, we construct the Trustworthiness Response Dataset (TRD) with 36,266 questions spanning four RAG settings. We reveal that existing approaches address isolated scenarios—prioritizing one knowledge source, naively merging both, or refusing answers—but lack a unified framework to handle different real-world conditions simultaneously. Therefore, we propose the BRIDGE framework, which dynamically determines a comprehensive response strategy of large language models (LLMs). BRIDGE leverages an adaptive weighting mechanism named soft bias to guide knowledge collection, followed by a Maximum Soft-bias Decision Tree to evaluate knowledge and select optimal response strategies (trust internal/external knowledge, or refuse). Experiments show BRIDGE outperforms baselines by 5–15% in accuracy while maintaining balanced performance across all scenarios. Our work provides an effective solution for LLMs’ trustworthy responses in real-world RAG applications.
1 Introduction
Recent advancements in retrieval-augmented generation (RAG) have enhanced the accuracy and trustworthiness of large language models (LLMs) in knowledge-intensive tasks, mitigating critical issues such as outdated model knowledge and hallucinations guu2020retrieval ; lewis2020retrieval . After retrieval, and before generating responses, LLMs often rely on two primary knowledge sources: context acquired through retrieval, referred to as external knowledge, and parametric knowledge stored within the LLM during training, termed internal knowledge. However, both knowledge sources exhibit limitations. External knowledge is prone to irrelevance, misinformation, or incompleteness due to corpus quality, retriever limitations, or query complexity chen2024benchmarking ; zou2024poisonedrag ; dai2024bias . For instance, in the NQ dataset kwiatkowski2019natural for open-domain question answering, wang2024astute employed Google Search as the retriever and searched the top 10 web pages for each question, finding that 34% of these pages were completely irrelevant to the answer. In the PopQA dataset mallen2022not , this percentage even reached 50%. Similarly, internal knowledge may be outdated or erroneous, as LLMs are constrained by training data noise and the rapid evolution of world knowledge weidinger2021ethical ; longpre2024pretrainer .
Figure 1: Four scenarios in real-world RAG.
In real-world RAG applications, the limitations of internal and external knowledge sources can lead to four distinct scenarios wang2024astute ; zhang2024evaluating (as illustrated in Figure 1). In the ideal case, LLMs should adopt scenario-specific strategies: (1) When both knowledge sources align and are correct, LLMs should be faithful to all knowledge to respond. (2) If external knowledge contains errors while internal knowledge remains reliable, LLMs should be faithful to internal knowledge. (3) Conversely, when internal knowledge is outdated or incorrect but retrieved context provides valid information, LLMs should anchor responses to external evidence. (4) When neither source contains sufficient or credible information, LLMs should refuse to answer.
Existing research addresses these scenarios separately. When facing scenario (2) or (3), some works advocate aligning LLM’s response with one of the knowledge sources (either internal or external knowledge) pan2023risk ; chen2024universal ; li2022large ; weller2022defending ; zhou2023context . Others attempt to integrate both sources shi2024trusting ; asai2023self ; xiang2024certifiably ; wang2024astute ; wei2025instructrag ; zhou2025trustrag but overlook scenario (4), failing to establish safeguards when both knowledge are invalid. A third line of work focuses exclusively on refusal mechanisms for exceeding the boundaries of knowledge sources cao2023learn ; song2025measuring , yet neglects knowledge integration strategies. In summary, these isolated approaches only use some specific strategies to handle one or two of these situations, failing to reveal the inadequacies of LLMs when faced with multiple RAG scenarios. Hence, there is also a lack of a synthesized framework capable of simultaneously addressing these diverse scenarios while ensuring LLMs’ trustworthiness in the RAG system.
To comprehensively evaluate LLMs’ response strategies under four RAG scenarios, we construct the Trustworthiness Response Dataset (TRD) containing 36,266 multiple-choice questions. Each question explicitly provides both external knowledge (simulating post-retrieval context) and internal knowledge (eliminating internal knowledge gaps between different LLMs). TRD comprises four RAG scenarios mirroring our proposed response strategies: (1) Faithful to All Knowledge (FA): We provide correct internal and external knowledge for the corresponding questions. (2) Faithful to Internal Knowledge (FI): We modify correct external knowledge with false facts to construct poisoned data. (3) Faithful to External Knowledge (FE): Standardizing data containing incorrect internal knowledge across LLMs from different families is complicated. Therefore, we use the latest Wikidata and Wikipedia facts (beyond the LLMs’ internal knowledge cutoff date) to construct questions and external knowledge. This ensures the models lack up-to-date internal knowledge, thereby forcing them to rely solely on external knowledge for answers. (4) Refuse to Answer (RA): While ensuring that internal knowledge is outdated, we introduce false facts to corrupt the correct external knowledge, rendering the questions unanswerable based on all knowledge sources. Our evaluation of TRD reveals that existing LLM families and RAG baselines fail to handle all scenarios effectively due to their lack of a unified understanding across these four cases.
However, in this task, human behavioral patterns offer novel insights for addressing these scenarios. For instance, when confident in our internal knowledge, we perform limited external searches just for verification fu2006suboptimal . Conversely, for uncertain or time-sensitive questions, we conduct extensive retrieval and evaluate consistency of external knowledge rieh2002judgment . When no credible evidence exists, humans readily acknowledge ignorance and refuse to answer the question gigerenzer2010rationality —a critical behavior that current RAG systems struggle to emulate. Inspired by these observations, we propose Biased Retrieval anD Generation Evaluation (BRIDGE), a framework that enhances the trustworthiness of LLMs in RAG systems. Unlike prior work that either enforces strict alignment to single knowledge sources (hard bias) or treats all sources equally (no bias), BRIDGE introduces soft bias: an adaptive weighting mechanism where an Allocator module determines whether to prioritize broad external knowledge retrieval (for consistency evaluation) or deeper internal knowledge generation (for fact verification). Based on the TRD training set data, Allocator supports both in-context learning (ICL) and group relative policy optimization (GRPO) paradigms. Then, based on the weighted knowledge, BRIDGE computes matching scores and evaluates consistency through a Maximum Soft-bias Decision Tree—a bias-guided decision tree enabling fine-grained response strategy selection (e.g., trusting internal knowledge, following external evidence, or refusing to answer). Additionally, it incorporates a reflection strategy to rectify decision errors during inference. In summary, our contributions are as follows:
  • To comprehensively evaluate LLM response strategies in real-world RAG, we construct TRD and use it to reveal the limitations of prior isolated solutions.
  • We propose the BRIDGE, a unified approach introducing soft bias to balance internal and external knowledge dynamically, a Maximum Soft-bias Decision Tree, and a Reflection module—covering all scenarios. Furthermore, BRIDGE can be applied to both ICL and GRPO settings.
  • Experiments show that BRIDGE demonstrates superior and balanced performance on TRD and other trustworthiness benchmarks for single RAG scenarios.
2 Related Works
Retrieval-augmented Generation. The paradigm of RAG has evolved significantly since its inception, driven by the need to build LLMs on external knowledge while mitigating hallucinations. Early foundational work established the core framework of RAG guu2020retrieval ; lewis2020retrieval ; borgeaud2022improving . Subsequently, its retrieval module has been optimized through techniques such as query rewriting zheng2023take ; dai2023promptagator ; chan2024rqrag , passage reranking glass2022re2g ; yu2024rankrag , and adaptive efficient retrieval strategies asai2023self ; jiang2023active ; yao2023react . In the generation module, recent efforts have focused on improving generation efficiency via prompt compression jiang2024piperag ; xu2024recomp ; cheng2024xrag and decoding optimization merth2024superposition ; jin2024ragcache ; wang2025speculative . These approaches primarily enhance individual RAG components in non-adversarial settings, whereas our research focuses on adversarial conditions (internal and external knowledge containing conflicts or noise).
Trustworthiness of LLMs in RAG. Ensuring the trustworthiness of LLMs in the RAG system has become a critical research area, particularly in adversarial environments. For scenarios where internal and external knowledge sources conflict, some studies align model responses with internal knowledge pan2023risk ; xu2023earth ; hong2024so ; weller2022defending , while others prioritize external knowledge li2022large ; gekhman2023trueteacher ; zhou2023context . Beyond these binary choices, recent efforts to improve the trustworthiness of LLM responses attempt to merge conflict sources through contrastive decoding jin2024tug ; shi2024trusting or integrating consistent portions of internal and external knowledge xiang2024certifiably ; wang2024astute ; wei2025instructrag ; zhou2025trustrag . However, these methods rely on carefully designed prompt templates to assess knowledge consistency. This matching strategy is sensitive to textual variations and lacks mechanisms for handling unanswerable questions. Previous work on enhancing LLM refusal capabilities in RAG systems relies on building expert knowledge bases cao2023learn or fine-tuning generative models song2025measuring . However, these methods often fail to integrate multiple knowledge sources and lack plug-and-play compatibility with black-box model-based RAG systems. Collectively, these works are orthogonal to each other, typically targeting only one or two scenarios. In contrast, BRIDGE provides a unified framework that systematically addresses four RAG scenarios through soft bias allocation and a scored decision tree.
3 TRD Construction
In this section, we explain the construction process of TRD. We formalize our task scenario as follows: given a natural language question $q$, the system retrieves external knowledge $K_{ext}$ and generates explicit internal knowledge $K_{int}$ by answering $q$ without accessing $K_{ext}$. Each question in TRD is associated with four options: one correct option, three incorrect options. One of the options is "I don’t know".. The dataset is divided into a training set (29,012 questions), a validation set (3,627 questions), and a test set (3,627 questions). TRD simulates four RAG scenarios displayed in Figure 1. Below, we detail the construction process.
3.1 Data Preparation
We utilize two primary datasets to help construct multiple-choice questions in TRD. (1) NQ: NQ is an open-domain question answering dataset released by Google in 2019 kwiatkowski2019natural . The accessible dataset contains 315,203 question-answer pairs (QA pairs) and provides a long answer (collected from Wikipedia by annotators) and a short answer (one or more entities) for each question. We use NQ to generate questions for the FA and FI categories. (2) TAQA: A dataset of 20,148 time-sensitive QA pairs derived from Wikipedia tables, where the answers to the questions change over time zhao2024set . We use TAQA to construct part questions for the FE and RA categories.
3.2 Question Generation
Faithful to All Knowledge. We select QA pairs from the NQ training set that contain long answers and short answers that are not timestamps. The short answer serves as the correct option. Using a popularity-based entity substitution method longpre2021entity , we generate two incorrect answers that have the same entity type as the short answer from Wikidata. The long answer is treated as external knowledge, while GPT-4-turbo generates internal knowledge based on the question and the short answer. We believe that explicitly providing internal knowledge helps to eliminate the parametric knowledge gap among different family LLMs during evaluation.
Faithful to Internal Knowledge. Following the same procedure as above, we select questions from the NQ validation set with long answers and non-temporal short answers, and generate two incorrect options of the same entity type using the popularity-based entity substitution method longpre2021entity . However, in this setting, we replace the correct entity in the long answer with one of the incorrect entities, resulting in misleading external knowledge. The internal knowledge, generated by GPT-4-turbo from the original short answer, remains correct.
Faithful to External Knowledge. To simulate scenarios where LLMs must rely solely on external knowledge due to outdated or absent internal knowledge, we define two distinct question sub-types: (1) Evolution: This category targets factual updates for existing facts/events (e.g., Who won the latest Nobel Prize in Literature?). LLMs trained with a cutoff date (February 2023 in this paper) cannot track such updates. The TAQA dataset focuses on this type of question. For each question, we query its latest answer in Wikipedia (February 1, 2025 dump). We retain only questions with valid Wikipedia pages. Then, using prompt templates in ConflictBank su2024conflictbank , we we use GPT-4-turbo to generate a Wikipedia-style evidence paragraph from the correct answer (post-cutoff date) as external knowledge. For the same question, we generate a parallel evidence paragraph describing TAQA’s 2021 answer (pre-cutoff date) using the same templates, simulating outdated internal knowledge. The correct answer is the entity with the post-cutoff date. We use the 2020 and 2021 answers in TAQA as incorrect options. (2) Perpetuation: This category addresses entirely new post-cutoff entities/events (e.g., What is the CPU of iPhone 16?). For many LLMs whose training data ends in 2023, it is unknown. First, we execute SPARQL [^1] https://www.wikidata.org/wiki/Wikidata:SPARQL_tutorial queries on Wikidata to extract timestamped triples where the object entity was created after February 2023 (e.g., <iPhone16, CPU, A18>). For these triples, we retrieve Wikipedia paragraphs (February 1, 2025 dump) containing both the subject and object entities as external knowledge; triples without corresponding paragraphs are filtered out. Internal knowledge is explicitly set to "None" to eliminate parametric knowledge interference from different LLMs. We select half of the data as FE data, and the other half is prepared to generate RA data.
Refuse to Answer. We replace the correct answer in external knowledge with an outdated distractor entity (e.g., altering A18 to A17 in the iPhone16 example). This ensures no credible evidence exists across all knowledge sources, asking LLMs to abstain.
3.3 Question Validation
We use GPT-4-turbo and Llama3 70B to validate questions, ensuring:(1) For FA, both models could select correct answers using either knowledge source. (2) For FI, models must choose the correct answer using internal knowledge but a consistent incorrect answer with external knowledge. (3) For FE, models must answer correctly with external knowledge but fail with internal knowledge. (4) For RA, both models select wrong answers using either knowledge source.
After validation, we balance the dataset, limiting FA to 10% of the total (since it is not the primary focus of our research) and allocating 30% each to the other three categories. Furthermore, the construction of time-related questions is automated, allowing them to be updated based on changes to Wikipedia and Wikidata, thus meeting the future evaluation needs of LLMs with evolving internal knowledge. For TRD construction details and statistics, please see Appendix A.
4 Biased Retrieval and Generation Evaluation
As illustrated in Figure 2, BRIDGE operates through two sequential phases: (1) Bias-Guided Knowledge Collection: This phase constructs an Allocator that computes retrieval dependency probability $r_{p}$ and generation dependency probability $g_{p}$ to determine the reliance degree on external and internal knowledge (constrained by $r_{p}+g_{p}=100\%$). We implement two Allocator paradigms: a reinforcement learning-based approach using GRPO and ICL method (§ 4.1.1). Subsequently, the system generates $n$ sub-queries capturing key information of $q$, and allocating the number of sub-queries based on $r_{p}$ and $g_{p}$. For these sub-queries, we yield generated knowledge $K_{gen}$ from Multi-query Generator and retrieved knowledge $K_{ret}$ from Multi-query Retriever (§ 4.1.2). (2) Bias-Guided Knowledge Evaluation: This phase employs a BGE-m3 chen2024bge based encoder as a Scorer to compute multi-granularity matching scores between knowledge sources (§ 4.2.1). These scores feed into a Maximum Soft-bias Decision Tree that determines the optimal response strategy under bias weights and thresholds. Furthermore, a Reflection module will trigger re-execution when the decision tree detects potential decision errors (§ 4.2.2).
Figure 2: Overview of BRIDGE.
4.1 Bias-Guided Knowledge Collection
4.1.1 Soft Bias Allocator
Determining soft bias (i.e., $r_{p}$ and $g_{p}$) for a single question presents significant challenges. For instance, the question What is the CPU of iPhone 16? contains no time-sensitive keywords, requiring LLMs to assign bias probabilities through analyzing question semantic and assessing internal knowledge boundaries. To enhance this capability, we leverage TRD’s training data to synthesize annotated examples with analysis paths. We develop two Allocator paradigms: (1) a GRPO-optimized Llama3-8B-Instruct model (pretrained cutoff: Feb 2023), and (2) an ICL-based approach.
Soft Bias Analysis Data Construction. We synthesize data through three steps: (1) We set the retrieval dependency probability $r_{p}^{hard}$ and the generation dependency probability $g_{p}^{hard}$ based on the scenario type in TRD, which can be regarded as the golden label of bias. For TRD’s FA/FI categories, we set $r_{p}^{hard}=0\%$, $g_{p}^{hard}=100\%$; for RA and FE, $g_{p}^{hard}=0\%$, $r_{p}^{hard}=100\%$. (2) Using DeepSeek-R1 and o3-mini as reasoning models, we prompt them to generate analysis path $a$ and $r_{p}^{soft}$ and $g_{p}^{soft}$ (10%-90%) under explicit directional constraints from hard bias labels (e.g., $g_{p}^{soft}>r_{p}^{soft}$ for FA/FI). (3) We select the generated output having a smaller root mean square loss between the soft bias and hard bias labels as high-quality data.
Allocator (GRPO). We employ GRPO to fine-tune our Allocator, using a composite reward function to optimize Llama3-8B-Instruct. We instruct the model to output a tuple $\gamma=(a_{pred},r_{p},g_{p})$, comprising the analysis path $a_{pred}$, predicted retrieval dependency probability $r_{p}$, and generation dependency probability $g_{p}$. The training incorporates four distinct reward components:
Direction Reward: Ensures primary direction alignment between the soft bias and its corresponding hard bias. For question $q$, the direction reward function is defined as:
$R(\gamma,q;direction)=\begin{cases}3&\mathrm{if~{}}\left(r_{p}>g_{p};r_{p}^{hard}>g_{p}^{hard}\right)\mathrm{~{}or~{}}\left(r_{p}<g_{p};r_{p}^{hard}<g_{p}^{hard}\right)\\ 0&\mathrm{otherwise}\end{cases}$ (1)
Format Reward: Verifies output format completeness of $\gamma$ through regular expression matching:
$R(\gamma,q;format)=\left\{\begin{array}[]{ll}1&\mathrm{if~{}}r_{p},g_{p}\mathrm{~{}and~{}}a_{pred}\mathrm{~{}in~{}}\gamma\\ 0&\mathrm{otherwise}\end{array}\right.$ (2)
Sum Reward: Enforces the $r_{p}+g_{p}=100\%$ constraint:
$R(\gamma,q;sum)=\left\{\begin{array}[]{cc}1&\mathrm{if~{}}r_{p}+g_{p}=100\%\\ 0&\mathrm{otherwise}\end{array}\right.$ (3)
Analysis Quality Reward: Promotes semantic similarity between the model’s predicted analysis $a_{pred}$ and high-quality analysis paths generated by larger reasoning models. We compute reward using BGE embeddings to measure textual similarity:
$R(\gamma,q;analysis)=BGE(a_{pred},a)$ (4)
Allocator (ICL). We also implement an in-context learning (ICL) paradigm by retrieving the top-k most similar training examples for each TRD. Instead of relying on soft bias outputs from larger reasoning models, we convert the gold hard-bias presets into soft-bias labels (e.g., 100%→90%, 0%→10%) to construct pseudo soft bias demonstrations. This ensures that BRIDGE’s performance improvements are not attributable to external reasoning model outputs. The analysis is discussed in § 5.2, while implementation details for GRPO and ICL in Allocator are provided in Appendix B.
4.1.2 Knowledge Collection
Relying solely on the original question for retrieval often yields limited information, prompting recent work to focus on improving retrieval performance in RAG systems to capture broader knowledge weller2022defending . Similarly, generating multiple responses for knowledge-related questions can improve LLM self-consistency and mitigate hallucinations chen2024universal . To expand the knowledge search space, we leverage LLMs to analyze a given question and identify the key information required for resolution. For each key information, the LLM generates a targeted query. These sub-queries act as "probes," enabling both the retriever and the LLM to explore a wider range of knowledge within their respective domains.
Specifically, we allocate sub-queries proportionally to each knowledge source based on the Allocator’s soft bias probabilities. Given $n$ total sub-queries, the number of queries assigned to retrieved knowledge $s_{r}$ and generated knowledge $s_{g}$ is computed as $s_{r}=\lceil r_{p}\times n\rceil$ and $s_{g}=\lceil g_{p}\times n\rceil$, respectively. The first $s_{r}$ sub-queries are processed by the Multi-query Retriever, aggregating external documents into the retrieved knowledge $K_{ret}$, while the first $s_{g}$ sub-queries are answered by the Multi-query Generator, forming the generated knowledge $K_{gen}$. This adaptive allocation improves the effectiveness of knowledge searching.
4.2 Bias-Guided Knowledge Evaluation
4.2.1 Knowledge Scorer
According chen2024bge , our BGE-based Scorer computes three similarity metrics between knowledge sources: sparse matching for lexical overlap, dense matching for semantic similarity, and ColBERT khattab2020colbert matching for token-level relevance, respectively weighted as 0.2, 0.4, and 0.4 by grid search in the TRD validation set. This multi-granular approach better assesses knowledge relevance.
4.2.2 Maximum Soft-bias Decision Tree
To determine the optimal integration strategy, we develop a Maximum Soft-bias Decision Tree that evaluates knowledge consistency through four knowledge matching scores and soft bias to help LLM make the final decision. As shown in Figure 2, we get four matching scores from Scorer. Based on these scores and soft bias, the tree executes three sequential checks:
Conflict Detection. Trigger FA if $S_{1}+S_{4}>(1-S_{1})+(1-S_{4})$ (implying high consistency in two knowledge sources), the LLM will integrate all knowledge.
Confidence Calculation. We compute the trustworthiness score for the retriever and the LLM:
$T_{Ret}=r_{p}[S_{3}+(1-S_{2})]$ (5)
$T_{LLM}=g_{p}[S_{2}+(1-S_{3})]$ (6)
This formulation reflects two key insights: (1) Higher dependency probability ($r_{p}$ or $g_{p}$) amplifies the impact of corresponding knowledge matches, and (2) A mismatch between knowledge from one source reduces the trustworthiness score, while a match between knowledge from another knowledge source also decreases its score. For example, poor alignment between generated knowledge and internal knowledge ($S_{2}\rightarrow 0$), along with good alignment between retrieved knowledge and external knowledge ($S_{3}\rightarrow 1$), both contribute to lowering $T_{LLM}$.
Threshold-based Decision. As shown in Figure 2, we employ two thresholds ($\alpha=0.5$, $\beta=1.1$) determined through grid search on TRD’s validation set. For cases where $T_{Ret}<T_{LLM}$, the system adopts FI when $T_{LLM}>\alpha$, otherwise triggering RA. When $T_{Ret}\geq T_{LLM}$, three outcomes emerge: (1) FE if $T_{Ret}\geq\beta$, (2) RA if $T_{Ret}<\alpha$, or (3) Reflection mechanism adjusts the sub-queries for intermediate cases ($\alpha\leq T_{Ret}<\beta$). After three unsuccessful cycles, the system defaults to RA as a conservative response. The asymmetric thresholds ($\beta>\alpha$) mean the retriever knowledge source requires higher trustworthiness ($\beta=1.1$) due to potential retrieval noise.
Finally, our framework dynamically adjusts response strategies based on predicted decisions: FA employs all knowledge; FI relies solely on $K_{int}$ and $K_{gen}$; FE prioritizes $K_{ext}$ and $K_{ret}$; and RA responds with "I don’t know". The theoretical analysis of GRPO, the hyper-parameter settings, and all system prompts in BRIDGE are detailed in Appendix B.
5 Experiments
5.1 Experimental Setups
Datasets. We evaluated all methods on four datasets. (1) TRD: Our constructed test set. (2) RealtimeQA (RQA) kasai2024realtime : A dynamic news QA platform. We select 468 questions emerging after October 2023, providing their corresponding evidence as external knowledge. This dataset aligns with the FE scenario. (3) HotpotQAPoisoned (HQAP) yang2018hotpotqa : We use the version from RobustRAG xiang2024certifiably , where external knowledge is corrupted via PoisonedRAG zou2024poisonedrag . Thus, we generate valid internal knowledge from correct answers using the method in § 3.2. This dataset corresponds to the FI scenario. (4) ConflictBank su2024conflictbank : A multiple-choice dataset containing temporal misalignment facts fabricated from Wikidata entries with future timestamps. We use the original correct evidence as internal knowledge and fabricated facts as external knowledge, and randomly sample 1,000 questions to evaluate LLMs under the FI scenario.
RAG settings & Baselines. (1) RAG settings: We evaluate RAG baselines using three LLMs (GPT-3.5-turbo, Qwen 72B, and Llama3-8B-Instruct) with parametric knowledge cutoffs before February 2023. To ensure fairness, all methods employ identical retriever parameters and are granted access to the English Wikipedia dump (cutoff date: February 1, 2025). (2) Baselines: Our baselines contain various aspects of LLM trustworthiness in RAG systems. Prompt-based methods include OPIN zhou2023context , USC chen2024universal , RobustRAG xiang2024certifiably , InstructRAGICL wei2025instructrag , AstuteRAG wang2024astute , and TrustRAG zhou2025trustrag . GPT does not support deterministic decoding; hence, we report results averaged from three runs. Fine-tuning-based methods encompass SelfRAG asai2023self , CAD shi2024trusting , InstructRAGFT wei2025instructrag and Trust-Align song2025measuring . Detailed settings are available in Appendix C.
5.2 Experiment Results
Main Results. As shown in Table 1, compared to vanilla RAG, OPIN, which prioritizes external knowledge, achieves notable gains on RQA, but suffers performance drops on HQAP and ConflictBank. Conversely, USC, which enforces internal knowledge consistency, excels on HQAP and ConflictBank but deteriorates sharply on RQA. Knowledge integration methods (RobustRAG, InstructRAG, AstuteRAG, and TrustRAG) demonstrate consistent improvements over vanilla RAG on TRD, underscoring the effectiveness of combining different knowledge sources. However, their lower RR scores indicate that their response strategies for refusing to answer are insufficiently considered.
CAD, which lacks mechanisms to resolve knowledge conflicts or reject uncertain questions, underperforms across most scenarios. In contrast, SelfRAG, InstructRAGFT and Trust-Align, which dynamically select knowledge sources or validate knowledge boundaries, achieve significantly higher RR rates. Nevertheless, their simple knowledge integration strategy limits broader applicability in all scenarios. As described in §4.1.1, BRIDGEICL, despite relying solely on templated soft bias demonstrations (without using data from large-scale reasoning models), outperforms baselines, validating benefits stem from soft bias design rather than synthetic data from larger reasoning models. BRIDGEGRPO further excels with its adaptive bias allocation, achieving balanced performance across all scenarios. Notably, BRIDGE has a sum of 100% for Acc and RR on the ConflictBank dataset when using GPT-3.5-Turbo and Qwen72B. This demonstrates our model’s highly reliable response strategy: when unable to provide correct answers using its internal knowledge, it consistently chooses to abstain from responding rather than risking incorrect outputs.
Table 1: Overall results, reporting accuracy (Acc), refusal rate (RR, regardless of question answerability), and exact match score (EM) rajpurkar2016squad , with the best results highlighted in bold.
Methods TRD RQA HQAP ConflictBank
Acc(%) RR(%) EM(%) EM(%) Acc(%) RR(%)
Based on GPT-3.5-turbo
Vanilla 53.71 5.81 72.50 44.40 52.04 0.20
OPIN zhou2023context 63.93 3.34 76.25 23.47 48.03 2.00
USC chen2024universal 65.53 3.01 26.25 70.70 73.46 1.08
RobustRAG xiang2024certifiably 62.50 6.69 65.82 64.00 58.40 3.04
InstructRAGICL wei2025instructrag 65.55 4.29 74.36 57.30 57.45 1.71
AstuteRAG wang2024astute 64.26 2.81 74.03 43.43 57.45 1.66
TrustRAG zhou2025trustrag 66.39 3.42 78.48 40.00 52.01 2.50
BRIDGEICL 72.32 31.40 76.92 70.97 61.86 38.14
BRIDGEGRPO 77.90 33.33 78.48 72.16 75.31 24.69
Based on Qwen 72B
Vanilla 63.19 9.50 70.25 61.61 57.83 0.61
OPIN zhou2023context 63.40 3.41 75.95 32.65 42.22 4.44
USC chen2024universal 64.20 4.80 28.75 76.00 70.74 2.04
RobustRAG xiang2024certifiably 64.33 7.28 74.25 65.00 59.72 3.04
InstructRAGICL wei2025instructrag 64.69 4.71 73.75 67.32 56.80 1.08
AstuteRAG wang2024astute 66.29 3.60 75.00 66.35 66.89 1.73
TrustRAG zhou2025trustrag 66.41 3.25 74.68 67.47 63.75 1.25
BRIDGEICL 74.20 39.56 76.25 72.15 67.35 32.65
BRIDGEGRPO 85.48 32.34 84.06 75.00 69.39 30.61
Based on Llama3 8B-Instruct
Vanilla 42.20 1.80 68.75 57.57 46.00 0.81
SelfRAG asai2023self 47.01 14.97 61.25 55.42 61.25 8.18
CAD shi2024trusting 24.40 1.50 55.00 35.00 39.20 1.95
InstructRAGFT wei2025instructrag 51.81 13.43 69.30 70.70 60.78 15.00
Trust-Align song2025measuring 53.05 18.11 72.50 69.30 65.61 5.10
BRIDGEICL 62.94 23.86 78.75 71.72 59.60 31.46
BRIDGEGRPO 73.33 19.30 80.00 78.72 71.72 28.28
Detailed Results. As shown in Figure 3, we compare the response accuracy of different methods across various RAG scenarios in TRD. BRIDGE demonstrates balanced capabilities across all scenarios, achieving performance comparable to state-of-the-art baselines on FA, FI, and FE cases while attaining optimal performance on RA scenarios. Figure 4 presents the error response strategy distribution of BRIDGE based on GPT-3.5-Turbo. The results reveal that FA exhibits the lowest error probability, whereas FI shows the highest. Notably, most FI cases are misclassified as FA, preserving the model’s ability to access the internal knowledge source. For FE misclassifications, the model typically adopts a conservative rejection strategy to maintain credibility. The misclassification of RA as FE represents an understandable scenario, as current RAG baselines tend to favor incorrect external knowledge over refusing to answer. Our method has minimized such potential harm through its carefully designed response mechanism.
Figure 3: Acc in different RAG scenarios.
Figure 4: Misclassification distribution of BRIDGE based on GPT-3.5-Turbo.
Table 2: Ablation study of BRIDGE, evaluated by Acc(%).
GPT-3.5-turbo Qwen 72B
Total FA FI FE RA Total FA FI FE RA
BRIDGEICL 72.32 87.93 92.48 61.64 56.58 74.20 82.98 92.17 54.62 74.78
BRIDGEGRPO 77.90 94.74 99.24 75.35 50.79 85.48 94.34 99.84 90.91 61.60
w/o Allocator 55.05 ($\downarrow$) 71.43 ($\downarrow$) 74.83 ($\downarrow$) 8.21 ($\downarrow$) 77.44 ($\uparrow$) 53.73 ($\downarrow$) 63.79 ($\downarrow$) 72.99 ($\downarrow$) 7.80 ($\downarrow$) 78.20 ($\uparrow$)
w/o Decision Tree (ICL) 66.67 ($\downarrow$) 94.83 ($\uparrow$) 97.74 ($\uparrow$) 79.31 ($\uparrow$) 7.75 ($\downarrow$) 69.12 ($\downarrow$) 82.84 ($\uparrow$) 97.35 ($\uparrow$) 41.66 ($\downarrow$) 13.81 ($\downarrow$)
w/o Decision Tree (GRPO) 66.74 ($\downarrow$) 98.25 ($\uparrow$) 96.21 ($\downarrow$) 78.87 ($\uparrow$) 7.94 ($\downarrow$) 72.22 ($\downarrow$) 90.57 ($\downarrow$) 98.10 ($\downarrow$) 82.29 ($\downarrow$) 19.32 ($\downarrow$)
Ablation Study. We perform an ablation study to evaluate the significance of Allocator and Decision Tree as shown in Table 2. (1) w/o Allocator: We remove the Allocator module, setting both $r_{p}$ and $g_{p}$ to 100% in BRIDGE. This eliminates soft bias guidance for knowledge collection and evaluation. The results show performance degradation in knowledge integration RAG scenarios (FA, FI, FE) within TRD compared to BRIDGE’s optimal performance.
Figure 5: Soft bias of Allocator for different RAG scenarios.
(2) w/o Decision Tree (ICL) & w/o Decision Tree (GRPO): Disabling the Maximum Soft-bias Decision Tree, we directly provide the LLM with knowledge and matching scores for final response generation. While this improved performance in some knowledge-integration scenarios, it caused significant performance drops in RA scenarios. The ablation study demonstrates that both modules play a critical role in enabling the LLM to maintain balanced response capabilities across all four RAG scenarios.
Allocator Performance. To further investigate the Allocator’s effectiveness, we examine its allocation behavior in BRIDGEGRPO mode across different RAG scenarios. As shown in Figure 5, the Allocator demonstrates accurate classification capability. The predominant allocation patterns of [80%-20%] and [20%-80%] indicate that the model can establish a correct soft bias direction for different RAG scenarios. This pronounced bias allocation significantly enhances the decision-making process within the decision tree, contributing to a reliable response.
Efficiency Analysis. In our experiments, the reflection mechanism is activated in approximately 10% of all cases. As shown in Figure 6, within these activated instances, the model’s problem-solving accuracy positively correlates with the number of reflection iterations. However, as a plug-and-play component, we carefully considered the efficiency trade-offs. Under ideal conditions, BRIDGE can resolve questions with 5 LLM API calls. When accounting for reflection operations, the average number of API calls per query reaches 5.89 on the TRD dataset - a number comparable to other prompt-based methods. Importantly, this modest increase in computational overhead enables BRIDGE to maintain balanced performance across diverse scenarios. We discuss the efficiency of the Reflection module in Appendix C.
Figure 6: Acc under different reflection iterations.
6 Conclusion
This paper addresses the critical challenge of balancing internal and external knowledge in retrieval-augmented generation (RAG) systems. First, we construct TRD to enable comprehensive evaluation. Then, we propose BRIDGE, a unified framework that introduces soft bias to dynamically weight knowledge sources, an Allocator module to guide retrieval/generation, and a Maximum Soft-bias Decision Tree to select optimal response strategies. Experiments demonstrate BRIDGE’s superiority over baselines, achieving balanced performance across all cases while significantly improving refusal rates in adversarial settings. Our work provides an effective and balanced approach to enhance the trustworthiness and robustness of RAG systems in real-world applications.
Appendix A Trustworthiness Response Dataset
A.1 Data Generation for Different RAG Scenarios.
A.1.1 FA Data
The NQ training set consists of 104,072 data samples. Following the method described in longpre2021entity , we perform entity matching based on popularity using the short-answer entities to generate two incorrect answers of the same type as the correct entity. We filter out invalid questions with empty long answers and samples where fewer than two corresponding entities could be retrieved, resulting in 76,409 valid QA pairs. For external knowledge, we use the provided long answers. For internal knowledge generation, referring to the prompt template of ConflictBank su2024conflictbank , we employ GPT-4-turbo with the following prompt template:
Prompt for Internal Knowledge Generation Based on the question and answer, create a piece of evidence that answers the question. 1. You cannot answer this question yourself. You must create evidence based on the answer I give. 2. This piece of evidence should be informative and well-structured. 3. The evidence should be presented in the tone of a Wikipedia entry. Question: {question} Answer: {correct_answer} Evidence:
A.1.2 FI Data
The NQ validation set contains 12,837 samples. Following a similar method to FA, we obtain 10,639 valid QA pairs after filtering and validation. We employ the approach from longpre2021entity to replace the long answer (external knowledge) with a false entity of the same type. For internal knowledge generation, we again utilize GPT-4-turbo with consistent prompt engineering to maintain comparability with the training setup.
A.1.3 FE Data
For the perpetuation data, we first identify entities emerging after February 2023 through SPARQL queries on Wikidata. The SPARQL query is:
1SELECT ?subject ?predicate ?object ?startTime WHERE {
2 ?subject ?date ?startTime .
3 ?subject ?predicate ?object.
4 FILTER(?startTime >= "2023-03-01T00:00:00Z"^^xsd:dateTime)
5}
We filter out meaningless attribute relations (e.g., subclass) to ensure data quality. Subsequently, we search Wikipedia for paragraphs containing both head and tail entities and set these paragraphs as external knowledge, discarding any triples without corresponding contextual passages. This process yields 9,523 valid triples with associated context. Using GPT-4-turbo, we convert these triples into question-answer pairs, where the tail entity serves as the answer, and the internal knowledge of all QA pairs is None.
For the evolution data, we leverage the TAQA dataset zhao2024set , which generates temporally-sensitive question-answer pairs from Wikipedia tables containing timestamp columns. TAQA’s methodology extracts event-related content from Wikipedia table rows at different timestamps, then employs GPT-4-turbo to generate corresponding questions and answers. We enhance this dataset through additional processing: while the original TAQA provides only tables as contextual knowledge, we utilize the WikipediaAPI to retrieve full articles corresponding to the most recent answers as external knowledge. For internal knowledge, we incorporate answers from the 2020 and 2021 versions. Finally, we get 11,326 QA pairs.
A.1.4 RA Data
We partition half of FE data to construct the RA data while preserving the original internal knowledge. The external knowledge is deliberately corrupted using the method from longpre2021entity to generate toxic context.
A.1.5 Data Validation Prompt
The verification prompt is as follows:
Prompt for Data Validation Answer the question by selecting the most accurate option based on the provided document; do not answer it by yourself. Return only the uppercase letter of the correct option. The output must follow this exact format: Correct Option: [Letter] Example: Correct Option: B Document: {internal_knowledge or external_knowledge} Question: {question} Options: {options}
A.2 Details and Statistics
To construct a focused evaluation dataset, we perform filtering to retain samples from the FI, FE, and RA categories. For the FA category, which is not our primary research focus, we apply random sampling to maintain its proportion at approximately 10% of the total dataset. Table LABEL:tab9 shows the data statistics,and Table LABEL:tab3-LABEL:tab8 shows some examples in TRD. The distribution of questions in TRD exhibits high similarity, significantly increasing the difficulty for the LLM to determine whether to rely on internal knowledge or external retrieval. For instance, questions such as who won first season of america’s got talent (Table LABEL:tab4) and who is the screenwriter of amelia’s children (Table LABEL:tab5) follow nearly identical syntactic patterns and lack explicit temporal keywords. This ambiguity necessitates careful knowledge calibration by the LLM to make appropriate source selection decisions.
Table 3: Statistics in TRD.
Total FA FI FE RA Avg $q$ Len Avg $K_{int}$ Len Avg $K_{ext}$ Len
Train 29,012 3,859 8,483 8,356 8,314 10.15 141.38 282.72
Val 3,627 486 1,071 1,056 1,014 10.18 143.54 284.42
Test 3,627 433 1,085 1,087 1,022 10.17 142.99 263.72
Total 36,266 4,778 10,639 10,499 10,350 10.16 142.64 276.95
Table 4: An example of FA.
Key Value
Question who played stefania ’s dad on everybody loves raymond
Question Type faithful to all knowledge
Temporal Fact Type none
Internal Knowledge David Proval’s Role as Stefania’s Dad on Ëverybody Loves RaymondD̈avid Proval is an accomplished American actor known for his versatile performances across television and film. In the popular sitcom Ëverybody Loves Raymond,ḧe portrayed the character of Frank, who is the father of Stefania, a recurring character introduced in the show’s later seasons. Stefania, played by actress Katia Winter, is the love interest of Robert Barone, one of the main characters. Proval’s portrayal of Stefania’s father added depth to the storyline, showcasing the dynamics of family relationships and cultural differences. His character was depicted as traditional and protective, often leading to humorous situations that highlighted the contrast between his values and those of the younger generation. David Proval’s performance was well-received, contributing to the show’s success and its ability to blend humor with relatable family themes. His role in Ëverybody Loves Raymondïs a testament to his talent and the show’s enduring popularity in American television history.
External Knowledge Marco Fogagnolo ( 6 episodes ; played by David Proval ) – The intimidating and threatening Italian father of Stefania , who takes over Nemo ’s Pizza after the family moves to America , jokingly implying that he “ did away with ” Nemo . Although he dislikes Robert , he befriends Frank and Marie but the two end up fighting for his friendship.
Internal Answer David Proval
External Answer David Proval
Options A. Fort Wayne B. I don’t know. C. Andrea D. David Proval
Correct Option D
Table 5: An example of FI.
Key Value
Question who won first season of americas got talent
Question Type faithful to internal knowledge
Temporal Fact Type none
Internal Knowledge Bianca Ryan - Winner of the First Season of America’s Got Talent. Bianca Ryan, a talented singer from Philadelphia, Pennsylvania, emerged as the winner of the inaugural season of "America’s Got Talent," which aired in 2006. At just 11 years old, she captivated both the judges and the audience with her powerful vocal performances, showcasing a maturity and skill beyond her years. Throughout the competition, Bianca delivered memorable performances, including renditions of "And I Am Telling You I’m Not Going" and "The Greatest Love of All." Her ability to connect emotionally with the audience and her impressive vocal range set her apart. In the finale, Bianca received a significant number of votes, securing her victory and a \$1 million prize. Her win established her as a prominent figure in the entertainment industry.
External Knowledge The first season of America’s Got Talent premiered on June 21, 2006 and concluded on August 17, 2006. The audition tour took place in April 2006, stopping at Los Angeles, New York, and Chicago. Regis Philbin hosted this season with judges David Hasselhoff, Brandy Norwood, and Piers Morgan. This season’s winner was Vicky Binns.
Internal Answer Bianca Ryan
External Answer Vicky Binns
Options A. Vicky Binns B. Lukas Graham C. I don’t know. D. Bianca Ryan
Correct Option D
Table 6: An example of FE (perpetuation).
Key Value
Question who is the screenwriter of amelia’s children
Question Type faithful to external knowledge
Temporal Fact Type perpetuation
Internal Knowledge none
External Knowledge Amelia’s Children is a 2024 Portuguese horror thriller film written and directed by Gabriel Abrantes and starring Jack Haven and Carloto Cotta.
Internal Answer none
External Answer Gabriel Abrantes
Options A.William Shakespeare B.Gabriel Abrantes C.Leonardo da Vinci D.I don’t know
Correct Option B
Table 7: An example of FE (evolution).
Key Value
Question which player is the most recent champion of the monterrey open singles
Question Type faithful to external knowledge
Temporal Fact Type evolution
Internal Knowledge Leylah Fernandez, a Canadian professional tennis player, emerged as the most recent champion of the Monterrey Open Singles, a prestigious event on the WTA Tour. The tournament, held annually in Monterrey, Mexico, attracts top talent from around the world and is known for its competitive field and vibrant atmosphere. In the latest edition of the Monterrey Open, which took place in March 2021, Fernandez showcased her exceptional skills and determination throughout the tournament. She advanced through the rounds with a series of impressive performances, ultimately reaching the final where she faced a formidable opponent. In a thrilling match that captivated tennis fans, Fernandez secured her victory, further solidifying her status as one of the rising stars in women’s tennis. This victory at the Monterrey Open marked a significant milestone in Fernandez’s career, adding to her growing list of achievements and enhancing her reputation on the international stage. Her triumph not only reflects her talent and hard work but also serves as an inspiration to aspiring athletes around the globe.
External Knowledge Donna Vekić, a professional tennis player from Croatia, is recognized as the most recent champion of the Monterrey Open Singles, a prestigious event on the WTA Tour. The tournament, held annually in Monterrey, Mexico, attracts top female tennis talent from around the world. Vekić secured her title at the 2023 edition of the tournament, showcasing her exceptional skills and determination on the court. In the final match, Vekić faced a formidable opponent, demonstrating her prowess through a combination of powerful serves and strategic gameplay. Her victory not only marked a significant milestone in her career but also highlighted her status as one of the rising stars in women’s tennis. The Monterrey Open, known for its vibrant atmosphere and enthusiastic crowds, provided an ideal backdrop for Vekić’s triumph, further solidifying her reputation in the sport. Vekić’s win at the Monterrey Open adds to her growing list of achievements and reflects her ongoing commitment to excellence in tennis. As she continues to compete at high levels, her performance in Monterrey serves as an inspiration to aspiring athletes and fans alike.
Internal Answer Leylah Fernandez
External Answer Donna Vekić
Options A. Donna Vekić B. Leylah Fernandez C. I don’t know. D. Leylah Fernandez
Correct Option A
Table 8: An example of RA (evolution).
Key Value
Question who is the most recent winner of the international brain bee
Question Type refuse to answer
Temporal Fact Type evolution
Internal Knowledge The International Brain Bee is a prestigious neuroscience competition that aims to inspire and encourage students to engage in the field of neuroscience. Held annually, the event attracts participants from around the globe, who compete in a series of challenges designed to test their knowledge and skills related to the brain and nervous system. As of the latest competition, the most recent winner of the International Brain Bee is Helene Li. Helene, representing her region, showcased exceptional knowledge in various areas of neuroscience, including neuroanatomy, neurophysiology, and the latest advancements in brain research. Her impressive performance during the competition not only earned her the title of International Brain Bee Champion but also highlighted her dedication to the study of neuroscience and her potential future contributions to the field. Helene Li’s achievement underscores the growing interest in neuroscience among young scholars and serves as an inspiration for future participants in the International Brain Bee, encouraging them to pursue their passions in understanding the complexities of the brain.
External Knowledge The International Brain Bee is a prestigious annual competition that challenges high school students from around the globe to demonstrate their knowledge of neuroscience and the brain. Established to promote interest in neuroscience, the competition includes a series of individual and team-based challenges such as quizzes, practical exams, and hands-on tasks related to brain function and neuroscience techniques. As of 2023, the most recent winner of the International Brain Bee is Rahil Patel. Representing [Insert Country or Region], Patel showcased exceptional knowledge and skills in the field of neuroscience, outperforming competitors from various international teams during the competition held in [Insert Location] on [Insert Date]. His victory not only highlights his individual achievement but also reflects the growing global interest in neuroscience among young minds and the importance of fostering scientific talent at an early age. Patel’s success exemplifies the opportunities provided by the International Brain Bee to inspire and educate future leaders in the field of neuroscience.
Internal Answer Helene Li
External Answer Rahil Patel
Options A. Rahil Patel B. Fredrick Odezugo C. I don’t know. D. Helene Li
Correct Option C
Table 9: An example of RA (perpetuation).
Key Value
Question in which country is stadionul dinamo located
Question Type refuse to answer
Temporal Fact Type perpetuation
Internal Knowledge none
External Knowledge In May 2021, Pablo Iglesias retired from politics after the 2021 Jerusalem regional election (in which he led the Podemos-IU list) delivered a resounding right-wing majority. As a result of his withdrawal, Iglesias returned to dedicate himself mainly to the media especially through his broadcast program by the Iranian channel HispanTV. In November 2022, Pablo Iglesias announced the start of a fundraising campaign for the launch of a leftist television channel, which would be called Canal Red, with the aim of competing with established channels that were considered conservative media by Iglesias and his followers,[3] the project had the backing of Jaume Roures, owner of Mediapro. The channel began its broadcasts through the internet on March 6, 2023.[1] In April 2023, the channel began broadcasting free-to-air television in Jerusalem through the frequency that 7NN, a conservative television channel that closed due to financial problems a month earlier, had occupied. The channel announced Inna Afinogenova as one of its first hires, Afinogenova is known in the Spanish-speaking world for having been one of the most visible faces of the Spanish version of RT, however, she resigned from the channel after the Russian invasion of Ukraine, despite that fact, some media continue to consider that she and consequently Canal Red are close to the influence of the Kremlin.
Internal Answer none
External Answer Jerusalem
Options A.I don’t know B.Caserta C.Jerusalem D.Madrid
Correct Option A
Appendix B The Details of BRIDGE
B.1 Allocator Based on GRPO
B.1.1 Data Generation
We use two reasoning models, DeepSeek-R1 and o3-mini, to generate analysis paths. The system prompt is as follows:
Prompt for Soft Bias Reasoning Data Generation For a given question, determine the required probability (10%-90%) of retrieving external knowledge versus answering directly based on stated human preference. Follow these steps: 1. ROLE DEFINITION: - Base knowledge cutoff: Strictly before Feb, 2023 - Never discuss human preference I provided in your response - Never disclose your knowledge cutoff date 2. TASK INSTRUCTIONS: Question: {question} Human Preference: {preference} (strong bias provided - MUST prioritize this direction) 3. PROCESSING STEPS: a) Analyze if the question requires: - Up-to-date information post-Feb, 2023 - Domain-specific expertise beyond general knowledge - Real-time/dynamic content (e.g., current events, live data) b) Analyze if the question can be answered using: - Pre-trained knowledge (before Feb, 2023) - Logical deduction - Common sense reasoning c) Strictly align probabilities with human’s specified preference direction while maintaining logical consistency 4. REQUIRED OUTPUT FORMAT: [Your Analysis] Probability of retrieving external knowledge: [10%-90%] (Needs external data: new information, specialized knowledge, or dynamic content) Probability of answering directly: [10%-90%] (Answerable using internal knowledge or reasoning.)
B.1.2 Theoretical Analysis of GRPO-based Allocator
The GRPO optimization process for the Allocator can be formalized through the following theoretical framework. Let $\pi_{\theta}(\mathfrak{a}|\mathfrak{s})$ denote the policy network parameterized by $\theta$, where $\mathfrak{s}=(q,K_{int})$ represents the state (question and internal knowledge) and $\mathfrak{a}=(r_{p},g_{p})$ denotes the allocation action (retrieval/generation probabilities).
Objective Function: The optimization objective combines four reward components:
$J(\theta)=\mathbb{E}_{(\mathfrak{s},\mathfrak{a})\sim\pi_{\theta}}\left[\sum_{i=1}^{4}\lambda_{i}R_{i}(\mathfrak{s},\mathfrak{a})\right]-\beta D_{KL}(\pi_{\theta}\|\pi_{ref})$ (7)
where $\lambda_{i}$ are reward weights, $R_{i}$ the reward functions, and $\beta$ controls KL-divergence from reference policy $\pi_{ref}$.
Policy Gradient: The gradient update rule follows:
$\nabla_{\theta}J(\theta)=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(\mathfrak{a}_{t}|\mathfrak{s}_{t})\left(\sum_{i=1}^{4}\lambda_{i}R_{i}-\beta\log\frac{\pi_{\theta}(\mathfrak{a}_{t}|\mathfrak{s}_{t})}{\pi_{ref}(\mathfrak{a}_{t}|\mathfrak{s}_{t})}\right)\right]$ (8)
Convergence Guarantee: Under Lipschitz continuity of reward functions ($\|R_{i}(\mathfrak{s},\mathfrak{a})-R_{i}(\mathfrak{s}^{\prime},\mathfrak{a}^{\prime})\|\leq L_{R}\|(\mathfrak{s},\mathfrak{a})-(\mathfrak{s}^{\prime},\mathfrak{a}^{\prime})\|$) and policy smoothness ($\|\pi_{\theta}-\pi_{\theta^{\prime}}\|\leq L_{\pi}\|\theta-\theta^{\prime}\|$), GRPO guarantees monotonic policy improvement with probability $1-\delta$ when:
$\beta\geq\frac{2L_{R}L_{\pi}\sqrt{|\mathcal{A}|}}{(1-\epsilon)^{2}}\sqrt{\log(1/\delta)}$ (9)
where $\epsilon$ is the discount factor and $|\mathcal{A}|$ the action space size.
This framework ensures the Allocator learns to make bias-aligned allocations while maintaining generation diversity through constrained policy optimization.
B.1.3 Training Details
We adopt the GRPO configuration for training parameters, specifically fine-tuning the Llama 3-8B-Instruct model on 8 Nvidia H100 GPUs with 80GB memory each. The key configurations include: employing LoRA hu2022lora adapters (rank=16, $\alpha$=32) to reduce memory requirements, using a learning rate of 5e-4 with a cosine scheduler, a per-device batch size of 32 combined with 4-step gradient accumulation, maximum prompt and generation lengths of 144 and 256 tokens respectively, and optimizing output quality through multiple reward objectives. For efficient distributed training, we implement data parallelism along with FlashAttention dao2022flashattention and bf16 mixed-precision training. All models are trained for 1 epoch using the AdamW loshchilov2017decoupled optimizer with a batch size of 32 by default. This setup ensures both computational efficiency and model performance while maintaining manageable memory usage.
B.2 Allocator Based on ICL
We employ BGE-m3 as the encoder to retrieve $k$ semantically similar demonstrations from the training set for each query. To determine the optimal value of $k$, we conduct hyper-parameter tuning on the validation set. A result is considered correct if the Allocator’s bias towards the question aligns with the hard bias. The experimental results, illustrated in Figure 7, demonstrate the impact of different $k$ values on model performance. We choose $k=5$; this setting enhances the model’s ability to leverage relevant contextual information while maintaining computational efficiency.
Figure 7: Hyper-parameter Learning for ICL.
B.3 Scorer Weight and Decision Tree Parameter Settings
We employ a grid search approach on the TRD validation set to optimize two critical sets of parameters: (1) the relative weights of the three-level matching scores in the Scorer module, and (2) the $\alpha$ and $\beta$ balancing thresholds in the Decision Tree module. Specifically, the relative weights are selected from the range $[0.1,1.0]$ with a step size of $0.1$. We also vary the $\alpha$ and $\beta$ balancing thresholds in $[0.1,2.0]$ with an interval of $0.1$. These carefully calibrated parameters, once determined, are maintained unchanged across all other datasets to ensure consistent evaluation standards.
B.4 All System Prompts in BRIDGE
For Allocator (GRPO) and Allocator (ICL), the prompt templates are as follows:
Allocator (GRPO) Task Description: Evaluate the following question to determine the probability of requiring external knowledge to answer it versus the probability of answering it directly [10%-90%]. Provide the results in the following format: Analysis: (Your analysis.) Probability of retrieving external knowledge: (Assess whether the question requires up-to-date data, specialized knowledge, or dynamic content.) Probability of answering directly: (Assess whether the question can be answered based on pre-trained knowledge.) Evaluate the following question: {question}
Allocator (ICL) Task Description: Evaluate the following question to determine the probability of requiring external knowledge to answer it versus the probability of answering it directly [10%-90%]. Provide the results in the following format: Probability of retrieving external knowledge: (Assess whether the question requires up-to-date data, specialized knowledge, or dynamic content.) Probability of answering directly: (Assess whether the question can be answered based on pre-trained knowledge or logical reasoning. Examples: {examples} Evaluate the following question: {question}
We fix the number of sub-queries $n$ to 10 across all experiments, a design choice that directly gets an integer when performing dependency probability multiplication operations.
Sub-query Generation Please design {number} new wildly diverse questions with different words that have the same answer as Original Question. Requirements: 1. Use different sentence structures. 2. Each question must employ a unique interrogative word (how/which/why, etc.). 3. Cover multiple dimensions of problem-solving. 4. Finally, rank the questions in descending order of importance for each dimension. Origin Question: {question} New Questions: 1. [New Question 1] 2. [New Question 2] {number}. [New Question {number}]
Multi-query Generator Please analyse before answering the following questions. If you are unsure of the answer or do not know the correct answer, please clearly respond with ’I don’t know’. Do not guess or make up information. Question: {generated_queries} Answers:
Our framework employs distinct response strategies based on the predicted decision: (1) For FA, the LLM generates responses using all available knowledge sources ($K_{int}$, $K_{gen}$, $K_{ext}$, $K_{ret}$) to provide comprehensive answers. (2) For FI, the LLM relies solely on internal knowledge ($K_{int}$) and generated knowledge ($K_{gen}$), ensuring factual consistency without external noise. (3) For FE, the response is derived from external knowledge ($K_{ext}$) and retrieved temporal knowledge ($K_{ret}$), prioritizing up-to-date information. (4) For RA, the LLM responds with "I don’t know" to avoid propagating potentially toxic or misleading content from adversarial external knowledge.
Responser Answer the question by selecting the most accurate option based on the provided document. Return only the uppercase letter of the correct option. The output must follow this exact format: Correct Option: [Letter] Example: Correct Option: B. Document: {knowledge} Question: {question} Options: {options}
Reflection Module ==Input data== Original question: {question} Knowledge document: Internal knowledge: {internal_knowledge} External knowledge: {external_knowledge} Generated knowledge: {generated_knowledge} Retrieved knowledge: {retrieved_knowledge} Phase 1: Knowledge contradiction analysis Please perform the following analysis steps: Consistency verification: 1. Confirm the consistency performance of internal knowledge and generated knowledge. 2. Confirm the consistency performance of external knowledge and retrieved knowledge. 3. Mark the specific contradictions between internal and external knowledge. Contradiction classification (analyzed from the following dimensions): Factual contradiction (objective fact difference) Timeliness contradiction (new and old information difference) Perspective contradiction (position/viewpoint difference) Integrity contradiction (information coverage difference) Root cause analysis: Model knowledge limitation (training data/time cutoff) External knowledge bias (source reliability/update frequency) Retrieval matching error (query-document relevance) Generate hallucination problem Phase 2: Problem reconstruction requirements Based on the above analysis, please design number new wildly diverse questions with different words that have the same answer as original question. Requirements: 1. Use different sentence structures. 2. Each question must employ a unique interrogative word (how/which/why, etc.). 3. Cover multiple dimensions of problem-solving. 4. Finally, rank the questions in descending order of importance for each dimension. 5. Pay attention to checking knowledge contradictions in the questions. Output format [Knowledge contradiction analysis] [Main contradiction] [Contradiction type] [Possible cause] [Reconstruct question list] (in descending order of importance) New questions: 1. [New Question 1] 2. [New Question 2] {number}. [New Question {number}]
B.5 Case Study
As shown in Figure 8, in the RA scenario, BRIDGE encounters a situation where the external knowledge is incorrect while the model’s internal knowledge remains empty. After retrieval and generation, the model fails to acquire valid knowledge, resulting in consistently low scores that lead to classification as a rejection scenario. In contrast, vanilla RAG methods are susceptible to interference from erroneous external knowledge, causing them to incorrectly select options like "B.Dream Works Animation" This demonstrates BRIDGE’s capability in handling knowledge-deficient scenarios through its rejection strategy.
Figure 8: A case study of BRIDGE.
Appendix C Baseline Settings
We segment the Wikipedia dump into passages of 256 tokens using sentence boundaries as delimiters, with no overlapping segments, and encode them using BGE-m3. This dump is contaminated with noisy data from the TRD dataset to simulate real-world scenarios involving unreliable external knowledge. For BRIDGE, the retriever defaults to fetching the single most relevant passage per sub-query. For other baselines, we allow retrieving up to 10 passages per question. This typically provides baselines with more retrieved knowledge than BRIDGE, representing a potential advantage. To ensure fair comparison, all models are given access to the same complete set of knowledge sources. Under ideal conditions, this setup ensures that their performance is not attributable to knowledge source availability constraints. We selected state-of-the-art representative baselines for evaluating trustworthiness in the RAG system. For the USC method, we permit 5 API calls to generate 5 responses while allowing integration of both internal and external knowledge sources. RobustRAG is configured to invoke an API for summarizing each retrieved passage, followed by keyword aggregation, resulting in over 10 API calls per question. For AstuteRAG and TrustRAG, whose performance typically correlates with knowledge integration iterations, we adopt the maximum values reported in their papers, ultimately requiring 4 API calls.
As shown in Table 10, we compare the API cost and efficiency between our method and existing approaches. To eliminate internal knowledge gaps across different LLM families, we directly generate the internal knowledge. Consequently, these generations are not counted toward the total API calls. As a result, BRIDGE requires an optimal of 4 API calls in the ideal case and a maximum of 7 in the worst case. While our method incurs slightly more API calls than some knowledge integration baselines, it achieves significantly better overall performance. We believe this represents a favorable trade-off, as the marginal increase in computational cost yields substantial gains in capability.
Table 10: API efficiency.
TRD ConflictBank
Methods Avg API Call Acc Efficiency Avg API Call Acc Efficiency
USC 6 65.53 10.87 6 73.46 12.24
RobustRAG 11 62.50 5.68 11 58.40 5.31
AstuteRAG 4 64.26 16.07 4 47.45 11.86
TrustRAG 4 66.39 16.60 4 30.00 7.50
BRIDGEICL 4.34 72.32 16.66 4.21 61.86 14.69
BRIDGEGRPO 4.22 77.90 18.46 4.14 78.12 18.87
Regarding fine-tuned approaches (SelfRAG, CAD, InstructRAGFT, and Trust-Align), we reproduce their training procedures by implementing the reported hyper-parameters and source code to train all models on Llama3-8B-Instruct. Notably, we adapt SelfRAG (originally implemented on Llama2-7B) following InstructRAGFT’s reproduction method, employing a learning rate of 1e-5 for 2 epochs. All models receive identical computational resources during training and inference phases.